OpenStudy (anonymous):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }+\frac{ 3*7 }{ 4*7}+\frac{ 3*7*11 }{ 4*7*10 }\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

the next one in line would be \[\frac{ 3*7*11*15 }{ 4*7*10*13 }\]

OpenStudy (anonymous):

i dont have a clue where to start here

OpenStudy (anonymous):

For what it's worth, you have the following closed form for the series: \[{\Large\sum_{n=1}^\infty}\frac{\displaystyle\prod_{k=0}^{n-1}(4k+3)}{\displaystyle\prod_{k=1}^n(3k+1)}\]

OpenStudy (anonymous):

Hmm, try the ratio test?

OpenStudy (anonymous):

thank you for the simplification there, although even simplified, it is confusing

OpenStudy (anonymous):

"Closed form" doesn't always mean "simplified" :P

OpenStudy (anonymous):

ok, i will have to look up closed form real quick

OpenStudy (freckles):

I feel like I seen a trick to this one but I can't remember off the top of my head.

OpenStudy (anonymous):

freckles!

OpenStudy (freckles):

Like on Openstudy recently

OpenStudy (anonymous):

i do believe it is the ratio test... it is soooo much tho

OpenStudy (anonymous):

wow, this one MUST be hard, if you two dont know :(

OpenStudy (anonymous):

Via the ratio test, \[\lim_{n\to\infty}\left|\frac{\displaystyle\prod_{k=0}^{n}(4k+3)}{\displaystyle\prod_{k=1}^{n+1}(3k+1)}\times\frac{\displaystyle\prod_{k=1}^n(3k+1)}{\displaystyle\prod_{k=0}^{n-1}(4k+3)}\right|=\lim_{n\to\infty}\left|\frac{4n+3}{3(n+1)+1}\right|\]

OpenStudy (anonymous):

ok,.... so, I need to learn how to do the closed form to find the initial values you got. then , is this \[\frac{ a_n+1 }{ a_n }\]?

OpenStudy (anonymous):

it looks like just a_n flipped over, thats y i was wondering

OpenStudy (anonymous):

You don't *need* the closed product formula for this one. The ratio test alone will show that a lot of lagging terms will cancel. And yes, the limit expression is \(\dfrac{a_{n+1}}{a_n}\).

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

Right, which means this series is...

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }>1\].... divergent?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

omg... idk how i would do this problem without learning what you did!

OpenStudy (freckles):

and i couldn't find any tricks i think it was just my mind playing tricks on me that there was some other trick or whatever anyways @SithsAndGiggles thingy worked out gorgeously

OpenStudy (anonymous):

for real! I thank you kindly!!!

OpenStudy (anonymous):

it is funny how it ended up being \[\frac{ 4n+3 }{ 3n+4 }\]

OpenStudy (anonymous):

Could either of you tell me if this could be considered a "taylor" series or something?

OpenStudy (anonymous):

i have several of these types of questions on my homework and I cant figure out how to do them. I don't even know what to call them. If you could lead me to something I could study myself, i would forever be grateful to you

OpenStudy (anonymous):

@SithsAndGiggles and @freckles

OpenStudy (anonymous):

You can find plenty of info and examples here: http://tutorial.math.lamar.edu/Classes/CalcII/SeriesIntro.aspx